# Bine Trees: Enhancing Collective Operations by Optimizing Communication Locality

### v1

@@ 3.2. Distance-Doubling Bine Trees
As with the distance-halving Bine tree, we first describe the case where the root is at rank 0. For a tree rooted at any other rank $t$, we follow the same approach as in distance-halving trees, by rotating rank identifiers to the right by $t$ positions.
@@ 3.2.1. Ranks Representation
We assign to each rank $r$ a value $h(r,p)$. For even ranks, h(r,p)=rank2nb(pr)h(r,p)=rank2nb(p-r), whereas for odd ranks $h(r,p)=rank2nb(r)$. Then, we assign to each rank $r$ a value $\nu(r,p)=h(r,p)\oplus(h(r,p)>>1)$ (>>>> is the logical right shift). Examples for even and odd ranks are shown in the right part of Fig. 6. When clear from the context, we omit $p$ for the sake of clarity.
@@ 3.2.2. Determining the Communication Partner
The communication partner of a rank $r$ at step $i$ is the rank $q$ such that $\nu(q)=\nu(r)\oplus 2^{i}$ (i.e., a rank that differs in the $i$-th bit in the $\nu$ representation). A rank communicates only after it receives the data from its parent. This happens at step $i$, where $i$ is the position of the most significant bit set to 11 in $\nu(r)$. In the example in the right part of Fig. 6, rank $2$ receives the data from its parent at step $1$, since the most significant bit set to 11 in $\nu(2)=011$ is in position $1$. Then, at step $2$, it sends the data to a rank whose $q$ such that $\nu(q)=011\oplus 2^{2}=011\oplus 100=111$ (i.e., rank $5$). The algorithm is correct because it operates in the same manner as the standard binomial tree algorithm, but using $\nu(r)$ instead of $r$.
@@ 3.2.3. Determining Ranks in a Subtree
All ranks in a subtree rooted at $r$ share the same $i+1$ least significant bits in their $\nu$ representation, where $i$ is the step at which $r$ received the data from its parent. Indeed, after receiving the data at step $i$, rank $r$ applies XOR on the bits in positions greater than $i$. This process recurses, and at each step the $i+1$ least significant bits remain unchanged. For example, in the right part of Fig. 6, rank $1$ receives the data at step $i=0$, and its $i+1=1$ least significant bit in $\nu(1)$ set. Consequently, all its descendants are ranks $q$ whose $\nu(q)$ also have the least significant bit set. It is worth noting that, unlike distance-halving Bine trees, nodes in distance-doubling Bine subtrees are not contiguous. For some collective operations, this may require transmitting non-contiguous data (see Sec. 4.3.1).
@@ 4.1. Gather
Similarly to standard binomial trees, in a Bine gather each rank gathers data from another rank at each step, halving the number of active ranks until only the root remains. We illustrate in Fig. 7, a standard binomial gather and a Bine gather over eight ranks, both using distance-halving trees. We consider data to be divided in “blocks”, one per rank participating in the gather. Let $[a,b]$ denote the range of blocks a rank holds at any time. When sending, a rank transmits all the blocks it currently has. In a standard binomial gather, if a rank holds $k$ blocks, it always receives the next $k$ contiguous blocks: that is, $[b+1,b+k]$. E.g., at step 1, rank 4 has blocks $[4,5]$ and receives $[6,7]$.
Figure 7. 8-node gather with binomial and Bine trees.
[…]
An 8-node gather with standard binomial trees and Bine trees.
@@ 4.3. Reduce-Scatter and Allgather
We discuss in this section the design of the reduce-scatter collective. For the allgather, it is enough to reverse the reduce-scatter communication pattern. We consider a reduce-scatter algorithm based on a vector-halving, distance-doubling butterfly pattern, using the Bine butterfly introduced in Sec. 3.1. In this scheme, each rank sends half of its data to a peer and receives half of the peer’s data at each step. Since the communication volume is highest in the early steps, relying on a distance-doubling butterfly helps reduce global traffic. As in standard recursive halving reduce-scatter, each rank sends a total of $n\frac{p-1}{p}$ bytes over $\log_{2}p$ steps.
A distance-doubling butterfly reduce-scatter can be conceptually viewed as the concurrent execution of a distance-doubling scatter and a distance-halving reduce. Focusing on the scatter component, unlike the scatter described in Sec. 4.2, a distance-doubling tree is used. While this reduces the volume of inter-group communication, it complicates data selection: a rank $r$ must send to a rank $q$ all blocks associated with the ranks in the subtree rooted at $q$. As discussed in Sec. 3.2.3, these subtrees consist of non-contiguous ranks, requiring the exchange of non-contiguous blocks.
@@ 4.3.1. Dealing with Non-Contiguous Data
Non-contiguous transmissions can hinder performance, especially for small- to medium-sized vectors (Loch and Koslovski, 2021). This can be mitigated in several ways. One option is to use MPI datatypes to handle non-contiguous data or to copy the data to a temporary buffer before transmission. However, this might still introduce a considerable overhead.
Alternatively, when $p$ is a power of two, we can permute the data upfront, ensuring subsequent transmissions use contiguous data. This can be achieved by moving the data blocks from position $r$ to position ${reverse}(\nu(r))$, where reverse{reverse} reverses the order of the bits in $\nu(r)$. As discussed in Sec. 3.2.3, all descendants of a given rank share a fixed number of least significant bits in their $\nu$ representation. By reversing $\nu(r)$, all blocks sent by rank $r$ would then share a fixed number of most significant bits, and therefore be contiguous. Another option is to skip the permutation and have ranks send contiguous data as if the permutation had been applied. For the reduce-scatter, this means that at the end of the collective each rank would hold the block of another rank. A final communication step is then sufficient to exchange data and reorder it correctly. Alternatively, some collectives combine two operations (e.g., an allreduce as a reduce-scatter followed by an allgather). In such cases, even if ranks do not receive the correct block after the first operation, the subsequent collective (e.g., the allgather) reverses the permutation and reorders the blocks correctly.
@@ 4.4. Allreduce and Alltoall
For the allreduce, we use two algorithms: one for small and one for large vectors. As for standard binomial tree algorithms, the former performs fewer steps, whereas the latter performs more steps but transfers less data. For small vectors, we employ a standard recursive doubling allreduce using Bine butterflies. For large vectors, we perform a Bine reduce-scatter followed by a Bine allgather. Both algorithms create a communication pattern similar to the Swing algorithm (De Sensi et al., 2024a). However, unlike Swing, which always sends non-contiguous data (Daniele De Sensi, 2025) and needs to issue multiple sends at each step, Bine sends contiguous data (when $p$ is a power of two), leading to $2\times$ performance improvements (see Sec. 5.2.2).
@@ 5. Experimental Evaluation
We implemented all Bine algorithms from Sec. 4 on top of MPI, without modifying any MPI library. This choice was motivated by both portability (our target systems rely on four different MPI libraries, see Table 2) and fairness, allowing for direct comparison against state-of-the-art algorithms using the same MPI backend. To validate this design choice, we implemented Bine versions of allreduce, allgather, and reduce-scatter within Open MPI and compared them to their counterparts built on top of MPI. The observed performance difference was below 3%. Therefore, in the following, we report results using the Bine implementations built on top of MPI. This was not the case on Fugaku, where, to concurrently use all the available NICs, our implementation uses the low-level uTofu library (Fujitsu, 2025) (see Sec. 5.4). We implemented all Bine algorithms in the PICO benchmarking framework (of Rome, 2025b), which streamlines comparison with the collective algorithms provided by default in Open MPI and MPICH. Additionally, we provide self-contained reference implementations (of Rome, 2025a).
@@
We compare Bine trees against all the collective algorithms available in the MPI implementations on each system. To ensure a fair comparison, we evaluate Bine both against the default algorithm selected by MPI and each available algorithm individually, by manually enforcing the selection. We also added other state-of-the-art algorithms from other MPI libraries or versions, such as the sparbit allgather algorithm (Loch and Koslovski, 2021; MPI, 2025d) (not available on Open MPI v4.1.x), and the ring-based allreduce algorithm (not available in MPICH 3.4a2, on which Cray MPICH v8.1.29 is based (MPICH, 2025)). Moreover, we implemented the Swing algorithm (De Sensi et al., 2024a) for allgather, reduce-scatter, and allreduce, and the torus-optimized Bucket algorithm (Jain and Sabharwal, 2010; Barnett et al., 1995) (essentially a multi-dimensional ring) for the same collectives. For each system, we begin with a focused comparison of Bine and binomial trees, and then extend the evaluation to include all available algorithms.
@@
For each collective, we measured performance across different node counts (without requesting any specific node placement) and vector sizes ranging from 32 B to 512 MiB. Due to space constraints, we report results only for power-of-two node counts. Results for non-power-of-two cases are consistent with those shown and are available in the reproducibility package. All collectives operate on vectors of 32-bit integers. For each combination of node count and vector size, we executed each collective up to $20000$ times for small vectors and five times for 512 MiB vectors. Most experiments were conducted with one process per node; performance with multiple processes per node is discussed in Sec.6. All vectors are located in host memory except for the experiments in Sec. 6 where we analyze multi-GPU collectives. We followed standard benchmarking practices (Hoefler and Belli, 2015), by discarding the first 20% of iterations as warm-up, and reporting for each iteration the maximum runtime across all ranks.
@@ 5.1. Evaluation on LUMI
LUMI (EuroHPC Joint Undertaking, [n. d.]) relies on a 24-group Slingshot 11 Dragonfly network (De Sensi et al., 2020), with 124 nodes per group. We used the LUMI-G partition, where each node has a 64-core AMD EPYC™ Trento 7A53 CPU (organized into four NUMA domains with 128GB of DDR4 memory each), four 200Gb/s Cassini-1 NICs, and four AMD MI250x GPUs.
@@ 5.1.1. Comparison with Binomial Trees
Given the large volume of data collected, across eight collectives, up to ten algorithms per collective, nine vector sizes, and up to seven node counts, for a total of over $5000$ configurations, we present a summarized analysis in Table 3444The complete set of detailed results is available in the reproducibility package.. For each collective, we report the fraction of configurations (i.e., combinations of node count and vector size) in which Bine outperforms binomial trees and vice versa (“%Win/%Loss”). When Bine is superior, we report both the average and maximum performance improvements over all the configurations. We compute the average using the geometric mean, which is more appropriate for summarizing ratios (Hoefler and Belli, 2015). Similarly, when binomial trees outperform Bine, we report the corresponding average and maximum performance degradations.
Lastly, we report both the average and maximum reduction in global-link traffic. Here, global links refer to those connecting switches in different Dragonfly groups. We assume packets traverse inter-group connections via minimal paths, an assumption that may not always hold in practice, especially in low-diameter networks such as Dragonfly (De Sensi et al., 2020). As a result, the reductions we report should be interpreted as lower bounds; the traffic savings induced by Bine trees could be even greater than our estimation.
@@
Bine reduces traffic on global links by an average of 10% across all collectives, and up to 20% at larger node counts, as more nodes lead to higher utilization of global links. The largest reduction, up to 94%, is observed in broadcast. This is due to MPICH (and Open MPI, as shown later) using a large-vector broadcast based on a distance-halving scatter, followed by a distance-doubling allgather (MPICH, 2025b; MPI, 2025c). In contrast, Bine trees reverse this process, significantly reducing traffic on global links, particularly in the allgather phase.
@@ 5.1.2. Comparison with Other State-of-the-Art Algorithms
Figure 8. Comparison with state-of-the-art algorithms on LUMI. (a): Each cell shows a letter indicating the best-performing algorithm (N = binomial, D = default), or, when Bine is best, the performance ratio between Bine and the next-best algorithm. (b): For each collective, the number below the name indicates the percentage of tests where Bine outperforms binomial trees. Box plots summarize the distribution of performance improvements in those cases.
[…]
Comparison of Bine performance against state-of-the-art algorithms on LUMI. Left: Each cell shows either a letter indicating the best-performing algorithm (N = binomial, D = default), or, when Bine is best, the performance ratio between Bine and the next-best algorithm. Right: For each collective, the number below the name indicates the percentage of configurations where Bine was the top performer. Box plots summarize the distribution of performance gains in those cases.
@@ 5.2. Evaluation on Leonardo
Leonardo (Turisini et al., 2024) uses a Dragonfly+ network topology based on Infiniband HDR. It consists of 23 groups, each containing 180 nodes, organized as two-level fat trees. We used the Booster partition, where each node is equipped with two 200 Gb/s dual-port NVIDIA Connect-X6 NICs, a single-socket 32-core Intel Xeon®8358 processor, 512 GB of host memory, and four NVIDIA A100 GPUs.

### v2

@@ 3.2. Distance-Doubling Bine Trees
We now describe the construction of distance-doubling Bine trees, using as a reference the one rooted in 0, shown with dashed blue lines in the right part of Fig. 6.
@@ 3.2.1. Ranks Representation
We assign to each rank $r$ a value h(r,p)=rank2nb(pr,p)h(r,p)=rank2nb(p-r,p) for $r$ even, and $h(r,p)=rank2nb(r,p)$ for $r$ odd. The only exception is $h(0,p)=0$. Then, we assign to each rank $r$ a value $\nu(r,p)=h(r,p)\oplus(h(r,p)>>1)$, with >>>> being the logical right shift. When clear from the context, we omit $p$.
@@ 3.2.2. Determining the Communication Partner
We describe the tree rooted at rank 0. For trees rooted at any other rank $t$, rank identifiers are rotated to the right by $t$ positions. The communication partner of a rank $r$ at step $j$ is the rank $q$ such that $\nu(q)=\nu(r)\oplus 2^{j}$ (i.e., a rank that differs in the $j$-th bit in the $\nu$ representation). A rank communicates only after it receives the data from its parent. This happens at step $i$, where $i$ is the position of the highest set bit in $\nu(r)$. In the example in the right part of Fig. 6, rank $2$ receives the data from its parent at step $1$, since the most significant bit set to 11 in $\nu(2)=011$ is in position $1$. Then, at step $2$, it sends the data to a rank whose $q$ such that $\nu(q)=011\oplus 2^{2}=011\oplus 100=111$ (i.e., rank $5$). In other words, the algorithm operates as the standard binomial tree algorithm, but using $\nu(r)$ instead of $r$, thus reducing the modular distance by $33\%$ (see Appendix A).
@@ 3.2.3. Determining Ranks in a Subtree
All ranks in a subtree rooted at $r$ have the same $i+1$ least significant bits in their $\nu$ representation, where $i$ is the step at which $r$ received the data from its parent. Indeed, after receiving the data at step $i$, rank $r$ applies XOR on the bits in positions greater than $i$. This process recurses, and at each step the $i+1$ least significant bits remain unchanged. E.g., rank $1$ receives data at step $i=0$, and its $i+1=1$ least significant bit in $\nu(1)$ is set. Consequently, all its descendants are ranks $q$ whose $\nu(q)$ also have the least significant bit set. Nodes in distance-doubling Bine subtrees are not contiguous, which may require transmitting non-contiguous data for some collectives (Sec. 4.3.1).
@@ 4.1. Gather
As in standard binomial trees, each rank gathers data from another rank at each step, halving the number of active ranks until only the root remains. We illustrate in Fig. 7, a binomial and a Bine gather over eight ranks, both using distance-halving trees. Data is divided into one “block” per rank. Let $[a,b]$ denote the range of blocks a rank holds at any time. When sending, a rank transmits all the blocks it currently has. In a standard binomial gather, if a rank holds $k$ blocks, it always receives the next $k$ blocks: that is, $[b+1,b+k]$. E.g., at step 1, rank 4 has blocks $[4,5]$ and receives $[6,7]$.
@@
Figure 7. 8-node gather with binomial and Bine trees.:
@@ 4.3. Reduce-Scatter and Allgather
We discuss in this section the design of the reduce-scatter collective. For the allgather, it is enough to reverse the reduce-scatter communication pattern. We consider a reduce-scatter algorithm based on a vector-halving, distance-doubling butterfly pattern, using the Bine butterfly introduced in Sec. 3.1. At each step, each rank exchanges half of its current data with a peer, so the size of the exchanged data halves at each step. Since the communication volume is highest in the early steps, relying on a distance-doubling butterfly helps reduce global traffic. As in standard recursive halving reduce-scatter, each rank sends a total of $n\frac{p-1}{p}$ bytes over $\log_{2}p$ steps.
A distance-doubling butterfly reduce-scatter can be conceptually viewed as the concurrent execution of a distance-doubling scatter and a distance-halving reduce. Focusing on the scatter component, unlike the scatter described in Sec. 4.2, a distance-doubling tree is used. While this reduces the volume of inter-group communication, it slightly complicates data selection: a rank $r$ must send to a rank $q$ all blocks associated with the ranks in the subtree rooted at $q$. As discussed in Sec. 3.2.3, these subtrees consist of non-contiguous ranks, requiring the exchange of non-contiguous data blocks, which can significantly hinder performance, especially for small- to medium-sized vectors (sparbit).
@@ 4.3.1. Dealing with Non-Contiguous Data
To mitigate the performance impact of transmitting non-contiguous data, we consider several possible options. One option is to use MPI datatypes to handle non-contiguous data or to explicitly copy the data to a temporary buffer before every transmission. However, this approach might still introduce a considerable overhead. Instead, for a reduce-scatter, we consider the following alternatives, which we compare in Appendix B:
Block-by-block:
Each block is transmitted independently. For small and medium-sized vectors, this approach increases the data transmission overhead. At the same time, it creates more opportunities for computation and communication overlap during reduction phases.
Permute:
Data is first permuted so that subsequent transmissions operate on contiguous regions of memory. This optimization is only applicable when $p$ is a power of two. The permutation is obtained by moving each data block from position $i$ to position ${reverse}(\nu(i))$, where reverse{reverse} denotes the bit-reversal of $\nu(i)$. As discussed in Sec. 3.2.3, all descendants of a given rank share a fixed number of least significant bits in their $\nu$ representation. By reversing $\nu(i)$, the transmitted blocks instead share a fixed number of most significant bits, making them contiguous in memory. For example, in Fig. 8, at step 0 of the reduce-scatter, rank 0 must send all blocks $i$ whose $\nu(i)$ representation has the least significant bit equal to 1 (see Sec. 3.2.3). That is, it must send blocks $1$, 22, 55, and $6$. After applying the permutation that places each block $i$ at position ${reverse}(\nu(i))$, these four blocks occupy positions 4477, enabling contiguous transmission.
Send:
Another option is to skip the permutation and let ranks transmit contiguous data as if the permutation had already been applied. In the example shown in Fig. 8, this means that, in the first step, rank 0 would send the last four blocks of the vector (i.e., $4$, 55, 66, and $7$ instead of $1$, 66, 22, and $5$). As a result, at the end of the collective, each rank would hold the block belonging to another rank. For instance, after the reduce-scatter, rank 1 would hold block $4$ (since block $1$ was not moved in position $4$). A final communication step can then be used to exchange and reorder the data correctly, with each rank $r$ sending its block to rank $q={reverse}(\nu(r))$. For example, rank 1 would send its block to rank $q={reverse}(\nu(1))=4$. It is worth noting that some collectives combine two operations (e.g., an allreduce implemented as a reduce-scatter followed by an allgather). In such cases, even if ranks do not hold the correct blocks after the first phase, the subsequent collective (e.g., the allgather) implicitly reverses the permutation and restores the correct order, thus avoiding the need for this additional communication step.
Two Transmissions:
A final option is to perform the reduce-scatter using a distance-halving rather than a distance-doubling butterfly. While this increases traffic over globally oversubscribed links, it simplifies handling non-contiguous data. Specifically, similar to the distance-halving scatter (Sec. 4.2), blocks are contiguous if the buffer is considered circular: any overflow at the end wraps around to the beginning. Thus, in practice this may require sending two segments: one at the end of the buffer and one at the beginning (for example, see rank 0’s buffer at step 1 of the Bine tree in Fig. 7).
Figure 8. Permutation of non-contiguous data blocks to guarantee contiguous transmissions.:
For an allgather, we have a symmetric scenario, i.e., in the permute case, the permutation is done at the end, after all the ranks received the entire buffer, and in the send case, the transmission to reorder the blocks is done before the actual steps of the collective operation.
@@ 4.4. Allreduce and Alltoall
For the allreduce, we use two algorithms: one for small and one for large vectors. As for standard binomial tree algorithms, the former performs fewer steps, whereas the latter performs more steps but transfers less data. For small vectors, we employ a standard recursive doubling allreduce using Bine butterflies. For large vectors, we perform a Bine reduce-scatter followed by a Bine allgather. Both algorithms create a communication pattern similar to the Swing algorithm (swing). However, unlike Swing, which always sends non-contiguous data (swing-git) and needs to issue multiple sends at each step, Bine sends contiguous data (when $p$ is a power of two), leading to $2\times$ performance improvements (see Sec. 5.2.2).
@@ 5. Experimental Evaluation
We implemented all Bine algorithms from Sec. 4 on top of MPI, without modifying any MPI library. This choice was motivated by both portability (our target systems rely on four different MPI libraries, see Table 2) and fairness, allowing for direct comparison against state-of-the-art algorithms using the same MPI backend. To validate this design choice, we implemented Bine versions of allreduce, allgather, and reduce-scatter within Open MPI and compared them to their counterparts built on top of MPI. The observed performance difference was below 3%. Therefore, in the following, we report results using the Bine implementations built on top of MPI. This was not the case on Fugaku, where, to concurrently use all the available NICs, our implementation uses the low-level uTofu library (utofu) (see Sec. 5.4). We implemented all Bine algorithms in the PICO benchmarking framework (pico; pico-github), which streamlines comparison with the collective algorithms provided by default in Open MPI and MPICH. Additionally, we provide self-contained reference implementations (bine-github).
@@
We compare Bine trees against all the collective algorithms available in the MPI implementations on each system. To ensure a fair comparison, we evaluate Bine both against the default algorithm selected by MPI and each available algorithm individually, by manually enforcing the selection. We also added other state-of-the-art algorithms from other MPI libraries or versions, such as the sparbit allgather algorithm (sparbit; ompi-allgather-sparbit) (not available on Open MPI v4.1.x), and the ring-based allreduce algorithm (not available in MPICH 3.4a2, on which Cray MPICH v8.1.29 is based (mpichnoring)). Moreover, we implemented the Swing algorithm (swing) for allgather, reduce-scatter, and allreduce, and the torus-optimized Bucket algorithm (bucket1; bucket2) (essentially a multi-dimensional ring) for the same collectives. For each system, we begin with a focused comparison of Bine and binomial trees, and then extend the evaluation to include all available algorithms.
@@
For each collective, we measured performance across different node counts (without requesting any specific node placement) and vector sizes ranging from 32 B to 512 MiB. Due to space constraints, we report results only for power-of-two node counts. Results for non-power-of-two cases are consistent with those shown and are available in the reproducibility package. All collectives operate on vectors of 32-bit integers. For each combination of node count and vector size, we executed each collective up to $20000$ times for small vectors and five times for 512 MiB vectors. Most experiments were conducted with one process per node; performance with multiple processes per node is discussed in Sec.6. All vectors are located in host memory except for the experiments in Sec. 6 where we analyze multi-GPU collectives. We followed standard benchmarking practices (htorbenchmarking), by discarding the first 20% of iterations as warm-up, and reporting for each iteration the maximum runtime across all ranks.
@@ 5.1. Evaluation on LUMI
LUMI (lumisupercomputerLUMIsFull) relies on a 24-group Slingshot 11 Dragonfly network (sc2020), with 124 nodes per group. We used the LUMI-G partition, where each node has a 64-core AMD EPYC™ Trento 7A53 CPU (organized into four NUMA domains with 128GB of DDR4 memory each), four 200Gb/s Cassini-1 NICs, and four AMD MI250x GPUs.
@@ 5.1.1. Comparison with Binomial Trees
Given the large volume of data collected, across eight collectives, up to ten algorithms per collective, nine vector sizes, and up to seven node counts, for a total of over $5000$ configurations, we present a summarized analysis in Table 3444The complete set of detailed results is available in the reproducibility package.. For each collective, we report the fraction of configurations (i.e., combinations of node count and vector size) in which Bine outperforms binomial trees and vice versa (“%Win/%Loss”). When Bine is superior, we report both the average and maximum performance improvements over all the configurations. We compute the average using the geometric mean, which is more appropriate for summarizing ratios (htorbenchmarking). Similarly, when binomial trees outperform Bine, we report the corresponding average and maximum performance degradations.
Lastly, we report both the average and maximum reduction in global-link traffic. Here, global links refer to those connecting switches in different Dragonfly groups. We assume packets traverse inter-group connections via minimal paths, an assumption that may not always hold in practice, especially in low-diameter networks such as Dragonfly (sc2020). As a result, the reductions we report should be interpreted as lower bounds; the traffic savings induced by Bine trees could be even greater than our estimation.
@@
Bine reduces traffic on global links by an average of 10% across all collectives, and up to 20% at larger node counts, as more nodes lead to higher utilization of global links. The largest reduction, up to 94%, is observed in broadcast. This is due to MPICH (and Open MPI, as shown later) using a large-vector broadcast based on a distance-halving scatter, followed by a distance-doubling allgather (mpich-bcast-sag; ompi-bcast-sag). In contrast, Bine trees reverse this process, significantly reducing traffic on global links, particularly in the allgather phase.
@@ 5.1.2. Comparison with Other State-of-the-Art Algorithms
Figure 9. Comparison with state-of-the-art algorithms on LUMI. (a): Each cell shows a letter indicating the best-performing algorithm (N = binomial, D = default), or, when Bine is best, the performance ratio between Bine and the next-best algorithm. (b): For each collective, the number below the name indicates the percentage of tests where Bine outperforms binomial trees. Box plots summarize the distribution of performance improvements in those cases.:
@@ 5.2. Evaluation on Leonardo
Leonardo (turisini2023leonardo) uses a Dragonfly+ network topology based on Infiniband HDR. It consists of 23 groups, each containing 180 nodes, organized as two-level fat trees. We used the Booster partition, where each node is equipped with two 200 Gb/s dual-port NVIDIA Connect-X6 NICs, a single-socket 32-core Intel Xeon®8358 processor, 512 GB of host memory, and four NVIDIA A100 GPUs.